### v1

@@ V Conclusion
We addressed urban 3D change detection in Subiaco using an object level, uncertainty aware framework on bi temporal LiDAR at city scale. The approach replaces perfect pre alignment assumptions with calibrated registration and a local level of detection, enforces class consistent instance association to resolve split and merge cases, and applies tiled processing with per object overlap, displacement, height, and volume cues to bound memory while preserving narrow ground changes. On 15 representative blocks the method reached 95.2% accuracy, 90.4% mF1, and 82.6% mIoU, improving over the strongest baseline by 0.2, 0.2, and 0.8 points, with a 7.6 point IoU gain on Decreased. The released Subiaco 2025 dataset and implementation enable reproducible evaluation and support HD map maintenance and smart mobility operations. Future work will extend to multi epoch sequences for trend analysis, fuse aerial and ground sensing, and learn thresholds that incorporate spatially varying uncertainty for online map updates.
Acknowledgment
This work was funded by the University of Jeddah, Jeddah, Saudi Arabia, under grant No.(UJ-24-SUTU-1290). The authors, therefore, thank the University of Jeddah for its technical and financial support. This project is a collaborative effort between the University of Jeddah and the University of Western Australia.

### v2

@@ V Conclusion
This paper presented a method for 3D change detection using an object level, uncertainty aware approach on bi-temporal LiDAR maps at city scale. Our approach replaces perfect pre-alignment assumptions with calibrated registration and a local level of detection. This enforces class consistent instance association to resolve split and merge cases. Our method applies tiled processing with per object overlap, displacement, height, and volume cues to bound memory while preserving narrow ground changes. We also proposed a new benchmark dataset to fill a significant gap in outdoor urban change detection research. Our method outperformed existing techniques on this challenging benchmark, highlighting the promise of our end-to-end workflow. The new Subiaco 2025 dataset and our implementation will enable reproducible evaluation and support HD map maintenance in addition to change detection. Future work will extend to multi epoch sequences for trend analysis, fuse aerial and ground sensing, and learn to incorporate spatially varying uncertainty for online map updates.
VI Acknowledgement
This work was funded by the University of Jeddah, Jeddah, Saudi Arabia, under grant No. (UJ-24-SUTU-1290). The authors, therefore, thank the University of Jeddah for its technical and financial support. This project is a collaborative effort between the University of Jeddah and the University of Western Australia, conducted under the Sustainable Partnerships Program with Distinguished International Universities.